### v1

@@ Abstract
Large Language Models (LLMs) have become integral to software engineering (SE), where they are increasingly used in development workflows. However, their widespread use raises concerns about the presence and propagation of toxic language—harmful or offensive content that can foster exclusionary environments. This paper provides a comprehensive review of recent research on toxicity detection and mitigation, focusing on both SE-specific and general-purpose datasets. We examine annotation and preprocessing techniques, assess detection methodologies, and evaluate mitigation strategies, particularly those leveraging LLMs. Additionally, we conduct an ablation study demonstrating the effectiveness of LLM-based rewriting for reducing toxicity. By synthesizing existing work and identifying open challenges, this review highlights key areas for future research to ensure the responsible deployment of LLMs in SE and beyond.
@@ 1 Introduction
Large Language Models (LLMs), such as BERT, GPT-4, Claude, and LLaMA, have seen extensive applications across a wide range of domains, including education, research, healthcare, software engineering (SE), and finance [16]. For instance, LLM-powered chatbots and virtual assistants are transforming customer support by offering efficient and personalized services [38]. In the healthcare sector, LLMs are contributing to advancements in medical research, drug discovery, and patient care [7]. As LLMs become more integrated into society, their widespread use has drawn increasing attention to their societal implications, both positive and negative.
@@
One pressing concern in the digital age is toxic language, which refers to harmful, offensive, or discriminatory language, including hate speech, personal attacks, threats, and harmful stereotypes. The prevalence of such toxic language in online platforms, such as social media and forums, creates hostile and exclusionary environments that hinder constructive dialogue and can cause significant harm to individuals [36]. Researchers are actively working to understand the underlying causes of toxic behavior and develop effective methods for its detection and mitigation. However, the field faces several challenges. Accurately identifying toxic language across diverse contexts and languages remains a significant obstacle. For example,  Radfar et al. [40] found that mildly offensive terms are more frequently used in tweets to express hostility between users with no social connection than between users who are mutual friends. Such complexities highlight the need for continued advancements in detection algorithms and comprehensive strategies to address toxicity.
@@
LLMs, with their ability to process and analyze vast amounts of textual data, offer promising solutions for addressing these challenges [13, 34]. Unlike traditional methods, LLMs can detect nuanced patterns of harmful language, including subtle and context-dependent forms of toxicity, that might otherwise go unnoticed. Researchers have also proposed and implemented various strategies to detect and mitigate toxicity in LLM-generated content, underscoring the potential of LLMs to play a central role in promoting healthier online interactions.
@@
This work synthesizes insights from a substantial body of publications on toxicity research conducted over the past four years. It delves into core topics such as toxicity detection and mitigation while also addressing critical auxiliary processes like corpus selection and data preprocessing. By consolidating findings from diverse studies, this paper provides a comprehensive overview of the current state of research and highlights gaps that require further exploration. One significant challenge with LLMs is their potential to generate toxic content, even in the absence of explicit prompts encouraging such behavior [20]. This issue stems from factors such as the presence of toxic content in training data [15], inherent algorithmic biases, or malicious exploitation. To better understand and address these issues, this paper proposes future research directions, including an ablation study that sheds light on the mechanisms underlying these behaviors and explores strategies for improvement.
@@
  • We systematically compare SE and non-SE datasets in the context of toxicity research
  • We categorize and summarize common annotation and preprocessing steps
  • We evaluate toxicity detection and mitigation solutions
  • We recommend using LLM-based rewriting approaches for toxicity mitigation, based on the results of our ablation study
The remainder of this paper is organized as follows: §2 introduces the use of LLM techniques in the SE domain, with a focus on toxicity analysis. §3 outlines common techniques for annotating, preprocessing, and utilizing datasets for toxicity detection and mitigation. §4 proposes improvements to pre-processing steps and toxicity detection. §5 details the setup, methodology, and results of our ablation study. The overall workflow is illustrated in Figure 1. Our code is available in an online repository: https://github.com/yangjack1998/SE-Mitigate
Figure 1: Overall Workflow of LLM Toxicity Detection and Mitigation
@@ 2 Background
This section presents the motivation and formal definition of toxicity studies within the SE community. Additionally, it highlights recent empirical works included in this review and outlines the criteria used for their selection.
@@ 2.1 Why is LLM in SE important
The integration of LLMs into SE has revolutionized various aspects of the field, offering substantial improvements in automation, efficiency, and decision-making. From code generation and summarization to commit classification and toxicity detection, LLMs are being increasingly leveraged to enhance software development workflows. This section explores the diverse applications of LLMs in SE, discussing how they contribute to key tasks such as classification and generation while also addressing their impact on developer interactions and the broader SE community.
2.1.1 Utilizing LLMs in SE
Software Engineering, a discipline that involves designing, developing, testing, and maintaining software, has hugely benefited from the advancement of LLMs. Many SE tasks can naturally be formulated as data analysis (learning) tasks such as classification tasks, regression tasks, and generative tasks [56]. These tasks are well-suited to being addressed by LLMs and hold significant potential for optimization and refinement through their application. For instance, automated code summarization, where the objective is to generate concise and descriptive natural language summaries for given code snippets or functions, can be naturally cast as a generative task.  Zhang et al. [59] propose a retrieval-based neural model to aid code summarization that improves the state-of-the-art methods by enhancing the model with the most similar code snippets retrieved from the training set.
@@ 2.1.1 Utilizing LLMs in SE
Hou et al. [25] point out that among their collection of LLM studies for SE tasks, following generation tasks, which the majority (about 70.97%) of studies center around, there are classification tasks that take up 21.62%. For example, categorizing commits into maintenance tasks helps practitioners make informed decisions regarding resource allocation and effectively manage technical debt. Researchers continue to explore ways to enhance the accuracy of this task through the application of LLMs. A study [21] proposed leveraging BERT (Bidirectional Encoder Representations from Transformers) to classify commits into three categories of maintenance tasks: corrective, perfective, and adaptive. Their approach demonstrated an approximately 8% improvement in accuracy compared to the previous state-of-the-art model.
2.1.2 Benefiting SE community using LLMs
@@
Toxic interaction can disrupt teamwork, discourage participation, and reduce productivity in software projects. Raman et al. [41] take an early step in implementing and illustrating a measurement instrument (an SVM classifier tailored for software engineering) to detect toxic discussions in GitHub issues. They also recognize the limited availability of large, software engineering-specific toxicity datasets, as well as the significant class imbalance they often exhibit (e.g., only 6 out of 1000 GitHub issues are toxic). To address this issue, Mishra and Chatterjee [34] explore the use of a zero-shot LLM (ChatGPT) for detecting toxicity in software-related text. Their findings show promising accuracy, underscoring the potential for integrating ChatGPT-enabled toxicity detection into developer communication channels. Currently, there is a lack of studies that go further to explore the mitigation of toxicity in software-related text using LLMs. To address this gap, a pilot study will be presented later in the paper to demonstrate the feasibility of such an approach.
@@
In terms of the distribution of LLM usages in SE activities, among six phases of the Software Development Life Cycle (SDLC) (i.e., requirements engineering, software design, software development, software quality assurance, software maintenance, and software management), over half of the studies are observed in the software development domain [25]. Within the domain, code generation and code completion are two tasks of great interest. While LLMs’ assistance with these tasks could largely improve the efficiency of SE practitioners, they have more subtle implications that we should be aware of. An empirical study [20] finds that pretrained neural language models (LMs) are susceptible to producing racist, sexist, or other forms of toxic language, posing challenges to their safe deployment. The tendency of pretrained LMs to generate toxic text even from seemingly harmless prompts highlights the importance of improving data selection processes during pretraining. This concern extends to similar generative processes, including code generation, as they are fundamentally driven by LLMs trained for text generation tasks.
@@ 2.2 What is toxicity
Understanding toxicity is crucial for fostering healthy interactions in online and professional communities, including software engineering. While the term ”toxicity” is broadly used to describe harmful or disruptive language, its precise definition and measurement can vary across contexts. This section explores how toxicity is defined in existing research, highlights the challenges in creating consistent annotation guidelines, and discusses key metrics used to evaluate both human annotations and model performance in toxicity detection.
2.2.1 Definition
@@
To study toxicity, defining the term is both essential and challenging, as it is broader and more subjective than other text classification problems, such as intent classification or spam detection [44]. The Perspective API [14] defines toxic content as ”texts that are rude, disrespectful, or unreasonable”. Gehman et al. [20] put more focus on the harm it creates, and describe toxic language as any form of speech or text that includes hate speech, offensive comments, harassment, or abusive behavior intended to harm, insult, or threaten others. A prior study [55] observed specific characteristics of toxicity definitions based on the domain of study. For example, in SE, the term is defined with more emphasis on its occurrence in conversations and distinctions such as ”offensive name-calling, insults, threats, personal attacks, flirtations, reference to sexual activities, and swearing or cursing” [46].
2.2.2 Metrics
@@
Annotation guidelines play a crucial role in ensuring consistency in toxicity detection datasets. Annotators follow standardized instructions to label instances of toxicity, helping to train models with reliable data. For example, the models powering certain toxicity detection systems, such as the Perspective API, rely on manual annotations guided by published annotation guidelines [2]. However, while clear and well-defined guidelines improve annotation accuracy and consistency, they do not eliminate subjectivity. Biases in guideline design, cultural differences among annotators, and varying interpretations of toxicity can introduce inconsistencies that affect the generalization of the model in different contexts [18]. A critical challenge in toxicity research is determining whether guidelines strike the right balance between recall and precision: Overly broad definitions risk flagging benign language, while narrow definitions may fail to capture implicit toxicity.
@@
To assess the reliability of these annotations, researchers measure inter-annotator agreement (IAA), a key metric that evaluates the consistency among annotators. Given the subjective nature of toxic language, ensuring high agreement is essential to produce reliable data sets [44, 46, 47]. High IAA indicates that annotation guidelines are effective and that annotators have a shared understanding of toxicity. Among the widely used methods for assessing IAA, Cohen’s $\kappa$ accounts for agreement occurring by chance and is calculated as $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$, where $P_{o}$ represents observed agreement and $P_{e}$ represents expected agreement by chance. Another commonly used metric, Krippendorff’s $\alpha$, generalizes inter-rater reliability by accommodating multiple annotators, diverse data types, and missing data. It is computed as $\alpha=1-\frac{D_{o}}{D_{e}}$, where $D_{o}$ represents observed disagreement and $D_{e}$ represents expected disagreement by chance. These measures are critical for validating the robustness and consistency of annotations in toxicity detection research. Yet, they have limitations—low agreement may indicate annotator disagreement, but it may also reveal deeper issues such as ambiguous labeling guidelines [42] or the need for more context-aware annotation strategies.
@@
Beyond annotation reliability, evaluating model performance is essential for assessing the effectiveness of toxicity detection systems. Performance metrics such as precision, recall, accuracy, F1 score, and F0.5 score provide insights into different aspects of model behavior. These metrics ensure a comprehensive understanding of how well a model balances false positives and false negatives. Table 1 presents the confusion matrix, which serves as the foundation for calculating these metrics, while Table 2 defines and details their respective equations. By employing rigorous evaluation metrics, researchers can gauge the strengths and limitations of toxicity detection models and refine them for improved real-world application.
Table 1: Confusion Matrix
@@
Table 2: Model Performance Metrics

### v2

@@ Abstract
Large Language Models (LLMs) have become integral to Software Engineering (SE), increasingly used in development workflows. However, their widespread adoption raises concerns about the presence and propagation of toxic language—harmful or offensive content that can foster exclusionary environments. This paper provides a comprehensive review of recent research (2020-2024) on toxicity detection and mitigation, focusing on both SE-specific and general-purpose datasets. We examine annotation and pre-processing techniques, assess detection methodologies, and evaluate mitigation strategies, particularly those leveraging LLMs. Additionally, we conduct an ablation study demonstrating the effectiveness of LLM-based rewriting for reducing toxicity. This review is limited to studies published within the specified timeframe and within the domain of toxicity in LLMs and SE; therefore, certain emerging methods or datasets beyond this period may fall outside its purview. By synthesizing existing work and identifying open challenges, this review highlights key areas for future research to ensure the responsible deployment of LLMs in SE and beyond.
@@ 1 Introduction
In Software Engineering (SE), toxic language within developer communications can hamper collaboration and productivity. Indeed, the increasing integration of Large Language Models (LLMs), such as BERT and GPT-4, into SE workflows, alongside their extensive applications across diverse domains like education, research, healthcare, and finance [dam2024complete], has amplified this concern. For instance, while LLM-powered chatbots and virtual assistants are transforming customer support by offering efficient and personalized services [pandya2023automating], and in the healthcare sector, LLMs are contributing to advancements in medical research, drug discovery, and patient care [athota2020chatbot], their widespread use has drawn increasing attention to their societal implications, both positive and negative.
@@
One pressing concern in the digital age is toxic language, which refers to harmful, offensive, or discriminatory language. The prevalence of such toxic language on online platforms, such as social media and forums, creates hostile and exclusionary environments that hinder constructive dialogue and can cause significant harm to individuals [mohan2017impact]. While commonly observed in general online discourse, toxicity in SE often manifests uniquely within technical contexts like code review discussions, issue trackers, and pull request comments, characterized by a blend of technical jargon and collaborative discourse. Researchers actively develop detection and mitigation methods, yet challenges persist in accurately identifying toxic language across diverse contexts and languages. For example, radfar2020characterizing found that mildly offensive terms are more frequently used in tweets to express hostility between users with no social connection than between users who are mutual friends.
@@
LLMs, with their ability to process and analyze vast amounts of textual data, offer promising solutions for addressing these challenges [cheriyan2021towards, mishra2024exploring]. Unlike traditional methods, LLMs can detect nuanced patterns of harmful language, including subtle and context-dependent forms of toxicity, that might otherwise go unnoticed. Researchers have also proposed and implemented various strategies to detect and mitigate toxicity in LLM-generated content, underscoring the potential of LLMs to play a central role in promoting healthier online interactions.
@@
This work synthesizes insights from a substantial body of publications on toxicity research conducted from 2020 to 2024. It delves into core topics such as toxicity detection and mitigation while also addressing critical auxiliary processes like corpus selection and data pre-processing. By consolidating findings from diverse studies, this paper provides a comprehensive overview of the current state of research and highlights gaps that require further exploration. One significant challenge with LLMs is their potential to generate toxic content, even in the absence of explicit prompts encouraging such behavior [gehman2020realtoxicityprompts]. This issue stems from factors such as the presence of toxic content in training data [gpt2knowyournumber], inherent algorithmic biases, or malicious exploitation. To better understand and address these issues, this paper proposes future research directions, including an ablation study that sheds light on the mechanisms underlying these behaviors and explores strategies for improvement.
@@
  • We systematically compare SE and non-SE datasets in the context of toxicity research.
  • We categorize and summarize common annotation and pre-processing steps.
  • We evaluate toxicity detection and mitigation solutions.
  • We recommend using LLM-based rewriting approaches for toxicity mitigation, based on the results of our ablation study.
While existing literature includes comprehensive surveys on offensive language and hate speech detection [mansur2023twitter, alkomah2022literature, jahan2023systematic], which primarily focus on general online discourse and algorithmic detection methods, more recent reviews have begun to specifically address toxicity in LLMs, encompassing both detection and mitigation strategies [villate2024systematic]. Our work, however, distinguishes itself through several key aspects:
  • Unique focus on toxicity in SE communications, addressing its specific contextual nuances.
  • Examination of challenges from LLM-generated toxicity and their mitigation role, including experiment-backed insights from an ablation study on LLM-based rewriting approaches—a novel contribution not found in existing surveys.
  • Detailed analysis of critical auxiliary processes: data annotation and pre-processing.
  • Consideration of ethical and policy implications for toxicity detection system deployment, an area less explored in existing reviews.
The remainder of this paper is organized as follows: §2 introduces the use of LLM techniques in the SE domain, with a focus on toxicity analysis. §3 outlines common techniques for annotating, pre-processing, and utilizing datasets for toxicity detection and mitigation. §4 outlines improvements in pre-processing and toxicity detection, alongside recommendations for ethical governance. §5 details the setup, methodology, and results of our ablation study. The overall workflow is illustrated in Figure 1. Our code is available in an online repository: https://github.com/yangjack1998/SE-Mitigate
Figure 1: Overall workflow of LLM toxicity detection and mitigation. Number annotations in each step represent relevant core literature as indexed in Table 5.
@@ 2.1 Why is LLM in SE important
The integration of LLMs into SE has revolutionized the field, offering substantial improvements in automation, efficiency, and decision-making. From code generation and summarization to commit classification and toxicity detection, LLMs are being increasingly leveraged to enhance software development workflows. This transformative impact stems from their ability to effectively process and understand both natural language descriptions and programming code, a capability exemplified by models like CodeBERT [feng2020codebert].
Many SE tasks can naturally be formulated as data analysis (learning) tasks such as classification tasks, regression tasks, and generative tasks [wang2022machine]. These tasks are well-suited to being addressed by LLMs and hold significant potential for optimization and refinement through their application. Specifically, the underlying transformer architectures [vaswani2023attentionneed] of modern LLMs, with their remarkable ability to model long-range dependencies and capture intricate contextual relationships within sequences, are particularly well-suited to the complexities of SE data.
For instance, automated code summarization, where the objective is to generate concise and descriptive natural language summaries for given code snippets or functions, can be naturally cast as a generative task.  zhang2020retrieval propose a retrieval-based neural model to aid code summarization that improves the state-of-the-art methods by enhancing the model with the most similar code snippets retrieved from the training set. Unlike traditional recurrent neural networks (RNNs) that can struggle with very long sequences common in code or extended developer discussions, transformers effectively process these lengthy contexts [vaswani2023attentionneed], making them highly advantageous for SE-specific tasks involving technical vocabulary and structured code elements.
@@
hou2024large point out that among their collection of LLM studies for SE tasks, following generation tasks, which the majority (about 70.97%) of studies center around, there are classification tasks that take up 21.62%. For example, categorizing commits into maintenance tasks helps practitioners make informed decisions regarding resource allocation and effectively manage technical debt. Researchers continue to explore ways to enhance the accuracy of this task through the application of LLMs. A study [ghadhab2021augmenting] proposed leveraging BERT (Bidirectional Encoder Representations from Transformers) to classify commits into three categories of maintenance tasks: corrective, perfective, and adaptive. Their approach demonstrated an approximately 8% improvement in accuracy compared to the previous state-of-the-art model.
@@
In addition to the direct automation of coding tasks, LLMs are also poised for benefiting the SE community itself by addressing social challenges. Toxic interaction can disrupt teamwork, discourage participation, and reduce productivity in software projects. raman2020stress take an early step in implementing and illustrating a measurement instrument (an SVM classifier tailored for software engineering) to detect toxic discussions in GitHub issues. They also recognize the limited availability of large, software engineering-specific toxicity datasets, as well as the significant class imbalance they often exhibit (e.g., only 6 out of 1000 GitHub issues are toxic). To address this issue, mishra2024exploring explore the use of a zero-shot LLM (ChatGPT) for detecting toxicity in software-related text. Their findings show promising accuracy, underscoring the potential for integrating ChatGPT-enabled toxicity detection into developer communication channels. Currently, there is a lack of studies that go further to explore the mitigation of toxicity in software-related text using LLMs. To address this gap, an ablation study will be presented later in the paper to demonstrate the feasibility of such an approach.
@@
In terms of the distribution of LLM usages in SE activities, among six phases of the Software Development Life Cycle (SDLC) (i.e., requirements engineering, software design, software development, software quality assurance, software maintenance, and software management), over half of the studies are observed in the software development domain [hou2024large]. Within the domain, code generation and code completion are two tasks of great interest. While LLMs’ assistance with these tasks could largely improve the efficiency of SE practitioners, they have more subtle implications that we should be aware of. An empirical study [gehman2020realtoxicityprompts] finds that pretrained neural language models (LMs) are susceptible to producing racist, sexist, or other forms of toxic language, posing challenges to their safe deployment. The tendency of pretrained LMs to generate toxic text even from seemingly harmless prompts highlights the importance of improving data selection processes during pretraining. This concern extends to similar generative processes, including code generation, as they are fundamentally driven by LLMs trained for text generation tasks.
@@ 2.2 What is toxicity
Understanding and defining toxicity is crucial for fostering healthy online and professional interactions within software engineering. Toxicity is a broad and more subjective challenge compared to problems like spam detection [sarker2022identification]. The Perspective API [conversationAI] defines toxic content as “texts that are rude, disrespectful, or unreasonable”.  gehman2020realtoxicityprompts put more focus on the harm it creates, and describe toxic language as any form of speech or text that includes hate speech, offensive comments, harassment, or abusive behavior intended to harm, insult, or threaten others. A prior study [villate2024systematic] observed specific characteristics of toxicity definitions based on the domain of study. For the purpose of this paper and in alignment with the professional nature of the SE domain, we adopt the comprehensive definition provided by sarker2020benchmark: “An SE conversation will be considered as toxic, if it includes any of the following: i) offensive name calling, ii) insults, iii) threats, iv) personal attacks, v) flirtations, vi) reference to sexual activities, and vii) swearing or cursing.”
@@
Annotation guidelines play a crucial role in ensuring consistency in toxicity detection datasets. Annotators follow standardized instructions to label instances of toxicity, helping to train models with reliable data. For example, the models powering certain toxicity detection systems, such as the Perspective API, rely on manual annotations guided by published annotation guidelines [annotation2018]. However, while clear and well-defined guidelines improve annotation accuracy and consistency, they do not eliminate subjectivity. Biases in guideline design, cultural differences among annotators, and varying interpretations of toxicity can introduce inconsistencies that affect the generalization of the model in different contexts [fairchild2020interrater]. A critical challenge in toxicity research is determining whether guidelines strike the right balance between recall and precision: Overly broad definitions risk flagging benign language, while narrow definitions may fail to capture implicit toxicity.
To address these challenges in practical settings, particularly within software engineering, detailed annotation rubrics are employed to guide human annotators. For instance, sarker2020benchmark empirically developed a rubric for classifying toxicity in SE conversations, which includes specific rules for identifying the following types of toxic content:
  • Profanity or curse words: Direct use of profane or vulgar language.
  • Acronyms referring to expletives: Initialisms or abbreviations that stand for expletives.
  • Insults towards a person or their work: Direct insults or demeaning remarks aimed at an individual or their professional contributions.
  • Identity attacks: Attacks targeting an individual’s identity, such as those based on race, gender, religion, or other protected characteristics.
  • Threats: Explicit or implicit threats of harm, intimidation, or violence.
  • References to sexual activities: Explicit or implicit mentions of sexual activities.
  • Flirtations: Unwelcome flirtatious comments or sexual advances.
Crucially, this rubric also distinguishes self-deprecating language (e.g., using words like “dumb” or “stupid” when referring to oneself or one’s own work) as non-toxic, recognizing its common use in SE communities for expressing mistakes without malicious intent. These rules provide concrete examples and decision-making criteria, helping annotators to accurately distinguish between technical disagreements and genuinely toxic interactions, thereby ensuring that collected data accurately reflects the nuances of toxicity within SE communications.
Despite comprehensive rubrics, the subjective nature of toxicity often leads to ambiguities in real-world contexts. Consider the following comment [raman2020stress]: “Let’s remove anti-cheat motion, which delivers more of a pain in the ass. After all, it is also possible to add a plugin.” While the phrase “pain in the ass” contains profanity that might be flagged as toxic by some guidelines, its use here could be interpreted by others as a colloquial, non-malicious expression of difficulty within a technical context. Such borderline cases highlight the inherent difficulties in achieving perfect consensus during annotation.
@@
Inter-annotator agreement (IAA) is a key metric to evaluate the consistency and reliability of annotations. Given the subjective nature of toxic language, ensuring high agreement is essential to produce reliable data sets [sarker2022identification, sarker2020benchmark, sarker2023toxispanse]. High IAA indicates that annotation guidelines are effective and that annotators have a shared understanding of toxicity. Among the widely used methods for assessing IAA, Cohen’s $\kappa$ [cohen1960coefficient] accounts for agreement occurring by chance and is calculated as $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$, where $P_{o}$ represents observed agreement and $P_{e}$ represents expected agreement by chance. In the “pain in the ass” example, if one annotator labels it as toxic ($A_{T}$) and another as non-toxic ($A_{N}T$), this disagreement directly impacts the observed agreement ($P_{0}$) and consequently lowers the Kappa score, signaling a need to refine guidelines or provide more context. Another commonly used metric, Krippendorff’s $\alpha$ [krippendorff2018content], generalizes inter-rater reliability by accommodating multiple annotators, diverse data types, and missing data. It is computed as $\alpha=1-\frac{D_{o}}{D_{e}}$, where $D_{o}$ represents observed disagreement and $D_{e}$ represents expected disagreement by chance. These measures are critical for validating the robustness and consistency of annotations in toxicity detection research. Yet, they have limitations—low agreement may indicate annotator disagreement, but it may also reveal deeper issues such as ambiguous labeling guidelines [sandri2023don] or the need for more context-aware annotation strategies.
@@
Model performance is another essential aspect used to assess the effectiveness of toxicity detection systems, beyond annotation reliability. Performance metrics such as precision, recall, accuracy, F1 score, and F0.5 score provide insights into different aspects of model behavior. These metrics ensure a comprehensive understanding of how well a model balances false positives and false negatives. Table 1 presents the confusion matrix, which serves as the foundation for calculating these metrics, while Table 2 defines and details their respective equations. By employing rigorous evaluation metrics, researchers can gauge the strengths and limitations of toxicity detection models and refine them for improved real-world application. These metrics are critically used in our ablation study (§5) to evaluate and compare different aspects of model performance.
Table 1: Confusion matrix in binary classification
@@
Table 2: Model performance metrics in binary classification